A rectangular lawn has dimensions 30m by 40m. A strip of uniform width is to be placed around the perimeter. How wide should the strip be so that exactly half the lawn gets mowed?
I tried (2x-30)(2x-40) and found x=5m but I was marked wrong by my teacher :(

Answer by scott8148(6628)

5m is correct, but what was your technique?

the lawn is 30 by 40 or 1200 sq m, so half would be 600 sq m

(30-2x)(40-2x)=600 ___ 1200-140x+4x^2=600 ___ 600-140x+4x^2=0

x^2-35x+150=0 ___ (x-30)(x-5)=0